# Robot-mediated physical Human-Human Interaction in Neurorehabilitation: a position paper

### v1

@@ VII-B What is the Role of Motivation, Engagement, and Gamification in RHHI?
Motivation plays a crucial role in NR, and RHHI has been recognized for its potential to enhance motivation. Increased motivation may lead to longer exercise durations and greater intensity, which in turn could improve functional outcomes. Several studies on RHHI have reported motivational benefits compared to exercising alone, likely due to the social interaction component [48, 37, 148, 74, 35].
Some research has further shown that competitive RHHI games can encourage longer and more active participation [38, 74], reinforcing the idea that social interaction enhances engagement. Additionally, Goršič et al. [149] found that replacing a human opponent with a human-like computer-controlled opponent resulted in lower motivation levels, suggesting that RHHI benefits may not be easily replicated by artificial agents [73]. However, gamification effects are not necessarily additive, as studies on virtual rehabilitation environments indicate that multiple gamification elements do not always enhance engagement [150].
Studying motivation in RHHI is challenging because gamification and social interaction are often examined separately from other factors (as discussed in Sec. VII-A). For instance, many gamification studies typically involve contactless games such as Pong or air hockey [37, 151, 38, 152, 74]. Additionally, social interaction and gamification may unintentionally alter task dynamics. For instance, human opponents behave more unpredictably than computer-controlled opponents, affecting motion dynamics. Similarly, if participants receive different interaction instructions [153], it may impact both social engagement and movement execution, making it difficult to isolate the specific mechanisms underlying changes in effort or movement quality.
@@ VII-C Who should be the partner in RHHI, and should they interact as opponents or teammates?
Patients in RHHI can interact with three potential groups of people: therapists, other patients, and healthy loved ones (e.g., spouses or friends). The most appropriate pairing depends on the interaction’s purpose, such as clinical assessment, treatment, or social engagement. For example, a clinical assessment scenario, as discussed in Baur et al. [54], requires a therapist as the partner. However, therapists may not be ideal partners for competitive games, as patients might not perceive them as fair opponents or could find clinical feedback distracting [37]. Even if another patient may seem like a better match for competition, studies suggest that intelligent difficulty adaptation algorithms can allow patients to compete with unimpaired loved ones despite motor ability differences [38]. This may be preferable due to the motivational benefits of exercising with loved ones. In fact, studies suggest that participants have a more positive experience when they “get along” with their partner [151, 37, 154]. Beyond partner type, individual personality traits and cognitive abilities must also be considered [154].
On the other hand, depending on partner type, interaction dynamics may need to be adjusted. As discussed in Sec. III, a virtual spring with adjustable stiffness can connect joint positions between users. While this method is typically used to link a patient with a therapist, some research suggests that novices (i.e., patients) may learn better when paired with other novices rather than experts [155]. Alternatively, if a virtual spring is activated intermittently, e.g., through an on/off button, even an unimpaired loved one could assist by selectively providing support at critical moments [148].
@@ VII-D How much do therapists rely on haptic feedback in RHHI?
Two primary interaction approaches in RHHI are discussed in the literature (Sec. III): unidirectional (one-way) and bidirectional (two-way). In the unidirectional approach [53, 60, 23], one user receives haptic feedback from their partner, while the partner does not perceive any interaction forces. This can be implemented by using IMU sensors to track movement [53], by recording and replaying executed trajectories [23, 60], or by setting a zero-stiffness spring in one robot. In this scenario, one user—likely the therapist—must rely solely on visual and auditory feedback for task execution.
In bidirectional interaction, both users receive haptic feedback while interfacing with separate robotic devices [85]. This enables a therapist to perceive their patient’s movements through force feedback, adding an additional communication channel beyond visual and auditory cues. While the unidirectional approach reduces costs by requiring only one robot, its effectiveness compared to bidirectional interaction remains unclear. Available studies have shown that unidirectional and bidirectional interaction during wrist and ankle trajectory tracking produce similar performance improvements for the less-skilled partner in pairs of healthy individuals [60]. However, it remains uncertain whether these findings hold when a therapist interacts with a patient with sensorimotor impairments. In clinical settings, therapists may compensate for the lack of haptic feedback by visually observing the patient’s performance and adapting their movements accordingly.
A hybrid approach, where haptic feedback varies based on motor skill differences, may offer a promising solution. In standard bidirectional RHHI, the more-skilled users must exert greater effort to compensate for the deviations of their less-skilled partner. This imbalance can make interactions physically demanding for the therapist when connected to a patient. Noccaro et al. [30] demonstrated that asymmetric connections with different stiffness for each partner can reduce the effort of the more-skilled user while balancing effort to match baseline conditions without haptic connection.
@@ VII-E Can RHHI scale NR while preserving individualized care for group therapy applications?
Some studies have shown that a single therapist can supervise multiple patients by using robotic devices in a rehabilitation setting [88, 156, 92, 93]. For instance, Miller et al. [92] proposed an automated robotic gym where patient-robot assignments dynamically adjust to optimize rehabilitation outcomes. While this study was conducted in simulation, several rehabilitation centers, such as Fondazione Don Carlo Gnocchi (Italy), Shirley Ryan AbilityLab (USA), and SRH Gesundheitszentrum Bad Wimpfen (Germany), have established robotic gyms equipped with multiple robots. These facilities have demonstrated outcomes comparable to traditional rehabilitation setups, with a single therapist effectively supervising three to four patients. However, none of these systems have yet implemented a networked RHHI framework that allows physical interaction between patients and therapists.
Integrating RHHI into robotic gyms could address several key challenges. Currently, therapists must divide their attention across multiple patients, increasing the risk of suboptimal therapy, patient fatigue, or, in worst-case scenarios, injuries if critical adjustments are missed. To mitigate these issues, therapists must gain a comprehensive understanding of each patient’s performance through multisensory channels, including visual observation, verbal communication, and haptic feedback. Additionally, centralized monitoring systems could aggregate patient data from all robots, reducing therapist workload and enhancing oversight. This framework could further benefit from dynamic adjustment systems [92, 93] which optimize patient-robot assignments and introduce patient-therapist or patient-patient coupling to maximize rehabilitation outcomes.
@@
VII-F How will future rehabilitation based on RHHI reshape the role of Therapists?
The evolution of robotics in rehabilitation follows trends seen in other fields, such as industrial automation. In the early phases of robotics, a strict separation existed between robots and humans. Indeed, industrial robots performed repetitive tasks in isolation from human operators for safety reasons. However, this division is narrowing, and the long-term goal of fully autonomous robots is gradually shifting towards collaborative systems [157].
Similarly, rehabilitation robotics is transitioning toward more interactive and adaptive systems, with RHHI enabling real-time therapist-patient interaction. Rather than replacing therapists, these technologies have the potential to enhance therapy outcomes while increasing the need for highly skilled professionals trained in advanced robotic-assisted rehabilitation. The benefits of RHHI have been widely discussed in this paper, their impact on therapist employment and job roles remains an important topic to be investigated.
Prior studies have reported therapists’ concerns about robots potentially replacing them in the future [22]. However, RHHI is designed to complement therapist expertise rather than substitute it. In this framework, robotic systems and therapist knowledge evolve together, mutually enhancing each other’s capabilities. While algorithms provide automation, they lack the ability to fully capture the nuanced decision-making and adaptability required for effective rehabilitation.
The integration of clinical expertise within the robotic control loop remains a key area for further research. Over the next years, studies should actively engage rehabilitation stakeholders—therapists, patients, clinicians, health service providers, and healthcare administrators—through qualitative research and feasibility testing within existing medical models. This will ensure that RHHI continues to enhance rehabilitation without compromising the essential human element of therapy.
@@ VII-G Is the impact of technologies needed for RHHI high enough to justify additional costs?
To rigorously evaluate the potential of RHHI, researchers should conduct larger clinical trials with comparisons to conventional manual therapy. So far, studies comparing robotic-assisted rehabilitation to traditional approaches have not produced results strong enough to justify the cost of robotic devices. To address the financial barriers that have limited widespread adoption, companies are now introducing Hardware-as-a-Service models [158], which offer robotic devices through leasing programs. This shift could significantly improve accessibility for clinics and researchers, enabling broader clinical validation of RHHI. As part of this effort, robust Health Technology Assessment will be essential to evaluate the clinical and economic value of RHHI solutions [159]. An important aspect to consider is the therapist’s willingness to adopt RHHI compared to conventional robotic approaches. The perception of greater control over the exercise and its impact on the patient may enhance perceived usability, potentially making RHHI a more effective and acceptable solution in clinical practice. To validate this hypothesis, it will be essential to include therapists in clinical trials through structured interviews and usability assessments—recognizing that in the RHHI framework, therapists are not merely facilitators but end-users alongside patients.
@@ VIII Conclusions
We advocate for a RHHI framework where robotic technology is not a substitute for therapists but rather an augmentation of their expertise and capabilities. While robots offer strength, precision, and measurable outcomes, therapists remain vital for delivering personalized, multisensory care and nuanced decision-making. Our multidisciplinary team—including clinicians, therapists, and engineers—has outlined key directions for future research. First and foremost, future efforts should focus on the design of robotic systems and control algorithms that prioritize task functionality, safety, and cost-effectiveness. In parallel, further development is needed to enable robust networking between devices, with particular attention to privacy and system stability. Moreover, RHHI integration into clinical practice requires the involvement of therapists when designing system features and defining key functional exercises. Fundamental research challenges were also identified and should be addressed to advance the field of RHHI. This requires deeper insight into motor learning mechanisms during RHHI, including motivation, engagement, partner characteristics, and interaction dynamics, supported by multidisciplinary research on social and physical interaction. A comprehensive understanding of all these aspects is essential to reveal the true impact of this technology and to assess its applicability in NR and the future role of the therapist.

### v2

@@ VII-B What is the role of motivation, engagement, and gamification in RHHI?
Motivation plays a crucial role in NR, and RHHI has been recognized for its potential to enhance motivation. Increased motivation may lead to longer exercise durations and greater intensity, which in turn could improve functional outcomes. Several studies on RHHI have reported motivational benefits compared to exercising alone, likely due to the social interaction component [GajadharDoubleFun, Gorsic2017CompetitiveIntensity, Baur2018RobotSupportedTraining, baur2023competitive, Mace2017BalancingTraining].
Some research has further shown that competitive RHHI games can encourage longer and more active participation [Gorsic2017AGame, baur2023competitive], reinforcing the idea that social interaction enhances engagement. Additionally, Goršič et al. [Gorsic2020EffectsGame] found that replacing a human opponent with a human-like computer-controlled opponent resulted in lower motivation levels, suggesting that RHHI benefits may not be easily replicated by artificial agents [takagi2017physically]. However, gamification effects are not necessarily additive, as studies on virtual rehabilitation environments indicate that multiple gamification elements do not always enhance engagement [laver2017virtual].
Studying motivation in RHHI is challenging because gamification and social interaction are often examined separately from other factors (as discussed in Sec. VII-A). For instance, many gamification studies typically involve contactless games such as Pong or air hockey [Gorsic2017CompetitiveIntensity, Novak2014IncreasingGameplay, Gorsic2017AGame, Pereira2019ImpactInvolvement, baur2023competitive]. Additionally, social interaction and gamification may unintentionally alter task dynamics. For instance, human opponents behave more unpredictably than computer-controlled opponents, affecting motion dynamics. Similarly, if participants receive different interaction instructions [takagi2023competition], it may impact both social engagement and movement execution, making it difficult to isolate the specific mechanisms underlying changes in effort or movement quality.
@@ VII-C Who should be the partner in RHHI, and should they interact as opponents or teammates?
Patients in RHHI can interact with three potential groups of people: therapists, other patients, and healthy loved ones (e.g., spouses or friends). The most appropriate pairing depends on the interaction’s purpose, such as clinical assessment, treatment, or social engagement. For example, a clinical assessment scenario, as discussed in Baur et al. [baur2019beam], requires a therapist as the partner. However, therapists may not be ideal partners for competitive games, as patients might not perceive them as fair opponents or could find clinical feedback distracting [Gorsic2017CompetitiveIntensity]. Although matching patients with others of similar ability may appear ideal for competitive settings, studies indicate that intelligent difficulty adaptation algorithms can enable individuals to engage in competition with unimpaired loved ones despite differences in motor ability [Gorsic2017AGame]. Such interactions may offer additional motivational benefits by allowing patients to exercise with familiar and emotionally supportive partners. In fact, studies suggest that participants have a more positive experience when they “get along” with their partner [Novak2014IncreasingGameplay, Gorsic2017CompetitiveIntensity, pereira2021use]. Beyond partner type, individual personality traits and cognitive abilities must also be considered [pereira2021use].
On the other hand, depending on partner type, interaction dynamics may need to be adjusted. As discussed in Sec. III, a virtual spring with adjustable stiffness can connect joint positions between users. While this method is typically used to link a patient with a therapist, some research suggests that novices (i.e., patients) may learn better when paired with other novices rather than experts [ganesh2014two]. Alternatively, if a virtual spring is activated intermittently, e.g., through an on/off button, even an unimpaired loved one could assist by selectively providing support at critical moments [Baur2018RobotSupportedTraining].
@@ VII-D How much do therapists rely on haptic feedback in RHHI?
Two primary interaction approaches in RHHI are discussed in the literature (Sec. III): unidirectional and bidirectional. In the unidirectional approach [amato2024unidirectional, short2025effects, Ganesh2014], one user receives haptic feedback from their partner, while the partner does not perceive any interaction forces. This can be implemented by using IMU sensors to track movement [amato2024unidirectional], by recording and replaying executed trajectories [Ganesh2014, short2025effects], or by setting a zero-stiffness spring in one robot. In this scenario, one user—likely the therapist—must rely solely on visual and auditory feedback for task execution.
In bidirectional interaction, both users receive haptic feedback while interfacing with separate robotic devices [koh2021exploiting]. This enables a therapist to perceive their patient’s movements through force feedback, adding an additional communication channel beyond visual and auditory cues. While the unidirectional approach reduces costs by requiring only one robot, its effectiveness compared to bidirectional interaction remains unclear. Available studies have shown that unidirectional and bidirectional interaction during wrist and ankle trajectory tracking produce similar performance improvements for the less-skilled partner in pairs of healthy individuals [short2025effects]. However, it remains uncertain whether these findings hold when a therapist interacts with a patient with sensorimotor impairments. In clinical settings, therapists may compensate for the lack of haptic feedback by visually observing the patient’s performance and adapting their movements accordingly.
A hybrid approach, where haptic feedback varies based on motor skill differences, may offer a promising solution. In standard bidirectional RHHI, the more-skilled users must exert greater effort to compensate for the deviations of their less-skilled partner. This imbalance can make interactions physically demanding for the therapist when connected to a patient. Noccaro et al. [noccaro2025robot] demonstrated that asymmetric connections with different stiffness for each partner can reduce the effort of the more-skilled user while balancing effort to match baseline conditions without haptic connection.
@@ VII-E Can RHHI scale NR while preserving individualized care for group therapy applications?
Some studies have shown that a single therapist can supervise multiple patients by using robotic devices in a rehabilitation setting [aprile2019improving, wuennemann2020dose, miller2022automated, adhikari2023learning]. For instance, Miller et al. [miller2022automated] proposed an automated robotic gym where patient-robot assignments dynamically adjust to optimize rehabilitation outcomes. While this study was conducted in simulation, several rehabilitation centers, such as Fondazione Don Carlo Gnocchi (Italy), Shirley Ryan AbilityLab (USA), and SRH Gesundheitszentrum Bad Wimpfen (Germany), have established robotic gyms equipped with multiple robots. These facilities have demonstrated outcomes comparable to traditional rehabilitation setups, with a single therapist effectively supervising two to four patients. However, none of these systems has yet implemented a networked RHHI framework that allows physical interaction between patients and therapists.
Integrating RHHI into robotic gyms could address several key challenges. Currently, therapists must divide their attention across multiple patients, increasing the risk of suboptimal therapy, patient fatigue, or, in worst-case scenarios, injuries if critical adjustments are missed. To mitigate these issues, therapists must gain a comprehensive understanding of each patient’s performance through multisensory channels, including visual observation, verbal communication, and haptic feedback. Additionally, centralized monitoring systems could aggregate patient data from all robots, reducing therapist workload and enhancing oversight. This framework could further benefit from dynamic adjustment systems [miller2022automated, adhikari2023learning] which optimize patient-robot assignments and introduce patient-therapist or patient-patient coupling to maximize rehabilitation outcomes.
@@
VII-F How will future rehabilitation based on RHHI reshape the role of therapists and caregivers?
The evolution of robotics in rehabilitation follows trends seen in other fields, such as industrial automation. In the early phases of robotics, a strict separation existed between robots and humans. Indeed, industrial robots performed repetitive tasks in isolation from human operators for safety reasons. However, this division is narrowing, and the long-term goal of fully autonomous robots is gradually shifting towards collaborative systems [nardo2020evolution].
Similarly, rehabilitation robotics is transitioning toward more interactive and adaptive systems, with RHHI enabling real-time therapist-patient interaction. Rather than replacing therapists, these technologies have the potential to enhance therapy outcomes while increasing the need for highly skilled professionals trained in advanced robotic-assisted rehabilitation. The benefits of RHHI have been widely discussed in this paper; their impact on therapist employment and job roles remains an important topic to be investigated to foster education and collaboration within an interdisciplinary community [motus_academy, polimi_rehabtech].
Prior studies have reported therapists’ concerns about robots potentially replacing them in the future [postol2023we]. However, RHHI is designed to complement therapist expertise rather than substitute it. In this framework, robotic systems and therapist knowledge evolve together, mutually enhancing each other’s capabilities. While algorithms provide automation, they lack the ability to fully capture the nuanced decision-making and adaptability required for effective rehabilitation. Beyond therapists, home rehabilitation with RHHI can actively involve caregivers by enabling them to guide patients through the robotic system. Currently, such guidance is often ad hoc and unstructured. RHHI could standardize this interaction, combining caregivers’ adaptability with the knowledge and capabilities of rehabilitation robots.
The integration of clinical expertise within the robotic control loop remains a key area for further research. Over the next years, studies should actively engage rehabilitation stakeholders—therapists, patients, caregivers, clinicians, health service providers, and healthcare administrators—through qualitative research and feasibility testing within existing medical models. This will ensure that RHHI continues to enhance rehabilitation without compromising the essential human element of therapy.
@@ VII-G Is the impact of technologies needed for RHHI high enough to justify additional costs?
To rigorously evaluate the potential of RHHI, researchers should conduct larger clinical trials with comparisons to conventional manual therapy. So far, studies comparing robotic-assisted rehabilitation to traditional approaches have not produced results strong enough to justify the cost of robotic devices. To address the financial barriers that have limited widespread adoption, companies are now introducing Hardware-as-a-Service models [forbes2025], which offer robotic devices through leasing programs. This shift could significantly improve accessibility for clinics and researchers, enabling broader clinical validation of RHHI. As part of this effort, robust Health Technology Assessment will be essential to evaluate the clinical and economic value of RHHI solutions [banta2003development]. An important aspect to consider is the therapist’s willingness to adopt RHHI compared to conventional robotic approaches. The perception of greater control over the exercise and its impact on the patient may enhance perceived usability, potentially making RHHI a more effective and acceptable solution in clinical practice. To validate this hypothesis, it will be essential to include therapists in clinical trials through structured interviews and usability assessments—recognizing that in the RHHI framework, therapists are not merely facilitators but end-users alongside patients.
@@ VIII Conclusions
We advocate for a RHHI framework where robotic technology is not a substitute for therapists but rather an augmentation of their expertise and capabilities. While robots offer strength, precision, and measurable outcomes, therapists remain vital for delivering personalized, multisensory care and nuanced decision-making. Our multidisciplinary team—including clinicians, therapists, and engineers—has outlined key directions for future research. First and foremost, future efforts should focus on the design of robotic systems and control algorithms that prioritize task functionality, safety, and cost-effectiveness. In parallel, further development is needed to enable robust networking between devices, with particular attention to privacy and system stability. Moreover, RHHI integration into clinical practice requires the involvement of therapists when designing system features and defining key functional exercises. Fundamental research challenges were also identified and should be addressed to advance the field of RHHI. This requires deeper insight into motor learning mechanisms during RHHI, including motivation, engagement, partner characteristics, and interaction dynamics, supported by multidisciplinary research on social and physical interaction.